OpenStudy (anonymous):

The population of Dos Pueblos High School increased from 1800 students ten years ago to 1926 students last year. What was the percent of the increase?

OpenStudy (anonymous):

1926-1800=126 \[\frac{126}{1800}=\frac{x}{100}\]126*100=x*1800 12600=1800x x = 12600/1800 x = 7 7% increase click god answer, if I deserve a medal,thank